OpenStudy (anonymous):

Calc help. (Will medal) Hello, I need to find the total distance this particle travled for x = 0 and x = 6. I integrated the velocity to find the postion. I pluged in 6 and 0 to the equation and got 35.2657. I do it on my calculator, I get 6.111. Please help.

OpenStudy (anonymous):

\[[\frac{ 2 }{ 3 }(2(x)+3)^{\frac{ 3 }{ 2 }}- \frac{ 2 }{ 3 }(x)^{2}]_{0}^{6}\]

OpenStudy (anonymous):

I typed it in my calculator right.

OpenStudy (anonymous):

The graph does go negative from [4,6]

OpenStudy (mathmale):

Where is the velocity equation that you've integrated?

OpenStudy (anonymous):

\[x \sqrt{2x+3} -3x dx\]

zepdrix (zepdrix):

Hmm it looks like it goes negative from 0 to 3, not from 4 to 6...

zepdrix (zepdrix):

Anyway, if the velocity function goes negative anyway, then it's telling you that the particle stopped and changed directions. So you would have to break up your integral at that point.

OpenStudy (anonymous):

Oh! I see now! So we have \[\int\limits_{0}^{3}v(x) and \int\limits_{3}^{6}v(x)?\]

OpenStudy (anonymous):

Sorry upper limit 4

zepdrix (zepdrix):

Since the velocity function is below the x-axis, the distance is going to come out negative (corresponding to moving backwards). So I guess we should throw some absolutes on that first integral,\[\large\rm \left|\int\limits_0^3v(t)dt\right|+\int\limits_3^6v(t)dt\]

OpenStudy (anonymous):

Ok! Thank you for your time!

zepdrix (zepdrix):

np